# Sums of square roots that are close to an integer

### 20 Jan 2024 v2

@@ 1.2. The sum of square roots problem
This is a special case of a notoriously difficult problem in Numerical Analysis and Computational Complexity Theory which has a number of different formulations. Following the formulation of Demaine, Mitchell and O’Rourke [7, Problem 33], the question is, for given values $n,k\in\mathbb{N}$, to understand the smallest positive value that can be attained by
@@
We return to our question: given $1\leq a_{1},\dots,a_{k}\leq n$, what is the smallest positive value of $\|\sqrt{a_{1}}+\dots+\sqrt{a_{k}}\|$? The cases $k=1,2$ have nice and satisfying solutions. However, already the case $k=3$ appears to be quite nontrivial. Generalizing the lower bounds $\gtrsim n^{-1/2}$ for $k=1$ and $\gtrsim n^{-3/2}$ for $k=2$ one might perhaps be inclined to believe that $\gtrsim n^{-5/2}$ is a natural guess for $k=3$. There is a very nice argument showing that it is at least $\gtrsim n^{-7/2}$ that was communicated to us by Arturas Dubickas and Roger Heath-Brown.
@@
Multiplying over all possible combinations of signs $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\in\left\{-1,1\right\}$,
$\prod_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\in\left\{-1,1\right\}}\left(\varepsilon_{1}\sqrt{a}+\varepsilon_{2}\sqrt{b}+\varepsilon_{3}\sqrt{c}\right)=(a^{2}+(b-c)^{2}-2a(b+c))^{2}\in\mathbb{N}.$
If $x=\sqrt{a}+\sqrt{b}+\sqrt{c}$ is not an integer, then this shows the existence of a real number $r\leq(3\sqrt{n})^{7}$ such that $x\cdot r\in\mathbb{N}$. This shows that $\|x\|\gtrsim 1/r\gtrsim n^{-7/2}$. ∎
As for upper bounds, we do not even know how to show the existence of $1\leq a,b,c\leq n$ such that $\|\sqrt{a}+\sqrt{b}+\sqrt{c}\|\ll n^{-3/2-\varepsilon}$. One would naturally expect many such numbers to exist and this seems to be a very interesting problem. There are some related problems: one that appears strictly easier concerns the set
$A=\left\{\sqrt{a}+\sqrt{b}\mod 1:1\leq a,b\leq n\right\}\subset[0,1]$
where it might be interesting to have a better understanding of its gap structure. We note that $A\cap[0,c/n^{3/2}]=\emptyset$ which follows from Angluin-Eisenstat argument. Is it true that $A$ intersects every interval $J\subset[0,1]$ with length $|J|\gtrsim n^{-3/2}$ Is it true that there are ‘few’ large gaps at scale $\sim n^{-3/2}$? This would be indicative of $\|\sqrt{a}+\sqrt{b}+x\|\lesssim n^{-3/2-\varepsilon}$ having a solution $1\leq a,b\leq n$ for ‘most’ $x\in\mathbb{R}$.
@@
The proof uses both the van der Corput method (needing $k\geq 241$) and the Vinogradov method: the implicit constant $c>0$ is very small. The argument presented in Proposition 1 generalizes to imply the lower bound $\gtrsim n^{1/2-2^{k-1}}$. A polynomial rate appears to be out of reach. One could ask whether this approach can lead to concrete estimates for small $k$. We present a simple example result in that direction.
Proposition 2.
There exist $\gtrsim n^{51/11}$ integers $1\leq a_{1},a_{2},\dots,a_{6}\leq n$ with
$0<\|\sqrt{a_{1}}+\sqrt{a_{2}}+\dots+\sqrt{a_{6}}\|\lesssim n^{-14/11}.$
Note that the exponent $14/11$ does not improve on the Angluin-Eisenstat bound $3/2$, however, it shows the existence of many such solutions. The argument follows from standard computations and the Huxley-Watt [13] exponent pair $(p,q)=(9/56,9/56)$. This estimate can surely be improved, however, it is perhaps less clear whether the exponent can be improved by very much (recall that for $k=6$ it is maybe close to $-5.5$). The exponent pair hypothesis, see §1.5, suggests that, conditional on very difficult problems being solved, one should expect the method to be able to attain a bound $\ll n^{-k/2+1/2}$ which is roughly half of what one expects.
@@ 1.5. Exponential sum estimates
We will argue that for a suitable choice of $s\ll n^{k/2}$, the exponential sums are sufficiently small to be absorbed by the main term which then guarantees the existence of (many) solutions. The main bottleneck appears to be getting good control on the exponential sum. It is not terribly difficult to show, for example using van der Corput’s method [25, Theorem 6.4], see also §2.4, that
$\left|\sum_{a=1}^{n}e^{2\pi i\ell\sqrt{a}}\right|\lesssim\ell$
which appears to be the optimal rate when $1\leq\ell\lesssim\sqrt{n}$. After that, things become less clear. The exponent pair hypothesis
@@ 2. Proof
2.2. Proof of Proposition 2
Proof.
We recall that it suffices to guarantee that
$\sum_{\ell=1}^{\infty}\widehat{h}(\ell)\left|\sum_{a=1}^{n}e^{2\pi i\ell\sqrt{a}}\right|^{k}\ll\frac{n^{k/2}}{s}.$
We use a special case of a rather big machinery, see [14, Chapter 8]. If a function $f:[N,2N]\rightarrow\mathbb{R}$ behaves like a monomial in the sense that
$|f^{j}(x)|\sim FN^{-j}$
and if $F\geq N$, then one expects estimates of the type
$\left|\sum_{N\leq x\leq 2N}e^{2\pi if(x)}\right|\lesssim_{\varepsilon}F^{p+\varepsilon}N^{q+1/2-p/2}.$
In our case, we have $f(x)=\ell\sqrt{n}$ and thus $F\sim\ell\sqrt{N}$. Choosing the explicit Huxley-Watt [13] exponent pair $(p,q)=(9/56,9/56)$, we have (ignoring losses at scale $\sim\ell^{\varepsilon}$ for ease of exposition)
$\left|\sum_{a=1}^{n}e^{2\pi i\ell\sqrt{a}}\right|\lesssim\begin{cases}\ell\qquad&\mbox{if}~{}\ell\lesssim\sqrt{n}\\ \ell^{9/56}n^{65/112}\qquad&\mbox{if}~{}\ell\gtrsim\sqrt{n}.\end{cases}$
Setting $s=n^{A}$ for some $A\geq 1/2$ to be determined, we have
$\sum_{\ell=1}^{\infty}\widehat{h}(\ell)\left|\sum_{a=1}^{n}e^{2\pi i\ell\sqrt{a}}\right|^{k}$ $\lesssim\sum_{1\leq\ell\leq\sqrt{n}}\frac{1}{s}\left|\sum_{a=1}^{n}e^{2\pi i\ell\sqrt{a}}\right|^{k}+\sum_{\sqrt{n}\leq\ell\leq s}\frac{1}{s}\left|\sum_{a=1}^{n}e^{2\pi i\ell\sqrt{a}}\right|^{k}$
$+\sum_{s\leq\ell\leq\infty}\frac{s}{\ell^{2}}\left|\sum_{a=1}^{n}e^{2\pi i\ell\sqrt{a}}\right|^{k}$
The first sum can be bounded from above by
$\sum_{1\leq\ell\leq\sqrt{n}}\frac{1}{s}\left|\sum_{a=1}^{n}e^{2\pi i\ell\sqrt{a}}\right|^{k}\lesssim\sum_{1\leq\ell\leq\sqrt{n}}\frac{\ell^{k}}{s}\lesssim n^{\frac{k+1}{2}-A}.$
When $k\geq 2$, we need not worry about the first sum since
$n^{\frac{k+1}{2}-A}\ll\frac{n^{k}}{n^{A}}=\frac{n^{k}}{s}.$
The second sum is bounded by
$\sum_{\sqrt{n}\leq\ell\leq s}\frac{1}{s}\left|\sum_{a=1}^{n}e^{2\pi i\ell\sqrt{a}}\right|^{k}\lesssim n^{65k/112-A}\sum_{\ell\leq n^{A}}\ell^{9k/56}$
which can be simplified to yield the bound
$\sum_{\sqrt{n}\leq\ell\leq s}\frac{s}{\ell^{2}}\left|\sum_{a=1}^{n}e^{2\pi i\ell\sqrt{a}}\right|^{k}\lesssim n^{65k/112+9Ak/56}.$
The third sum, via a similar computation, is bounded by
$\sum_{s\leq\ell\leq\infty}\frac{s}{\ell^{2}}\left|\sum_{a=1}^{n}e^{2\pi i\ell\sqrt{a}}\right|^{k}$ $\lesssim n^{65k/112+A}\sum_{s\leq\ell\leq\infty}\frac{\ell^{9k/56}}{\ell^{2}}.$
In order for the infinite sum to converge, we require $k\leq 6$. Assuming $k\leq 6$, we get
$\sum_{s\leq\ell\leq\infty}\frac{s}{\ell^{2}}\left|\sum_{a=1}^{n}e^{2\pi i\ell\sqrt{a}}\right|^{k}\leq n^{65k/112+9Ak/56}$
which coincides with the second sum. In order to have these sums be $\ll n^{k}/s=n^{k-A}$ we require
$\frac{65k}{112}+\frac{9Ak}{56}<k-A$
which leads to
$A<\frac{47k}{18k+112}.$
Setting $k=6$ implies the Proposition. ∎
We note that the proof is lossy in many different ways: different exponent pairs may lead better results in different summation regimes and it is clear that some improvements are possible. It is less clear whether substantial improvements (say ones that beat the $n^{-3/2}$ bound) are possible.
@@ 2.2. Proof of Proposition 2
One of the big advantages is that the exponential sum $\sum_{1\leq a\leq n}e^{2\pi i\ell\sqrt{a}}$ is completely explicit. One way one could hope to avoid it is to use van der Corput’s approximation [25, Theorem 6.4] to get
@@
Ignoring the constants, there are two terms. One is harmless
$\lesssim\frac{1}{\nu}\qquad\mbox{and summing}\qquad\sum_{\ell/\sqrt{n}\leq\nu\leq\ell}\frac{1}{\nu}\lesssim\log n.$
The other term is, up to constants, given by
$\exp\left(\frac{i\pi\ell^{2}}{2\nu}\right)\frac{\ell}{\nu^{3/2}}\text{erfi}\left(\frac{\left(\frac{1}{2}+\frac{i}{2}\right)\sqrt{\pi}\left(\ell-2\nu\sqrt{x}\right)}{\sqrt{\nu}}\right)\big{|}_{1}^{n}$
Using
$\forall~{}x\in\mathbb{R}\qquad\left|\text{erfi}\left(\left(\frac{1}{2}+\frac{i}{2}\right)x\right)\right|\leq 2,$
one could ignore the term completely and would end up with
$\sum_{\ell/\sqrt{n}\leq\nu\leq\ell}\frac{\ell}{\nu^{3/2}}\lesssim\ell^{1/2}n^{1/4}$
which corresponds to the classical exponent pair $(1/2,0)$. It also implies, trivially,
$\sum_{\ell/\sqrt{n}\leq\nu\leq\ell}\frac{\ell}{\nu^{3/2}}\lesssim\ell\sum_{\ell/\sqrt{n}\leq\nu\leq\ell}\frac{1}{\nu^{3/2}}\lesssim\ell$
which appears to be sharp in the regime $\ell\leq\sqrt{n}$. However, the presence of a completely explicit term that induces further oscillations does seem to inspire some hope that are more precise understanding might be possible.

### 04 Mar 2024 v6

@@ 1.2. The sum of square roots problem
This is a special case of a notoriously difficult problem in Numerical Analysis and Computational Complexity Theory which has a number of different formulations. Following the formulation of Demaine, Mitchell and O’Rourke [9, Problem 33], the question is, for given values $n,k\in\mathbb{N}$, to understand the smallest positive value that can be attained by
@@
We return to our question: given $1\leq a_{1},\dots,a_{k}\leq n$, what is the smallest positive value of $\|\sqrt{a_{1}}+\dots+\sqrt{a_{k}}\|$? The cases $k=1,2$ have nice and satisfying solutions. However, already the case $k=3$ appears to be quite nontrivial. There is a very nice argument showing that it is at least $\gtrsim n^{-7/2}$ that was communicated to us by Arturas Dubickas and Roger Heath-Brown.
@@
Multiplying over all possible combinations of signs $\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\in\left\{-1,1\right\}$, one can define a polynomial
$p(x)=\prod_{\varepsilon_{1},\varepsilon_{2},\varepsilon_{3}\in\left\{-1,1\right\}}\left(x-(\varepsilon_{1}\sqrt{a}+\varepsilon_{2}\sqrt{b}+\varepsilon_{3}\sqrt{c})\right).$
One sees, by direct computation, that it is a polynomial all of whose coefficients are integers. It is also a classical result, see Besicovitch [3, 4], that if $a,b,c$ have different square-free parts, then $1,\sqrt{a},\sqrt{b},\sqrt{c}$ are linearly independent over $\mathbb{Q}$. This means that if $n\in\mathbb{N}$, then $p(n)\neq 0$ and thus $|p(n)|\geq 1$. Suppose now that $\sqrt{a}+\sqrt{b}+\sqrt{c}$ is not an integer but very close to the integer $n$. Using $|p(n)|\geq 1$ and that the product of the other 7 factors being $\leq(3\sqrt{n})^{7}$ now shows the result. ∎
As for upper bounds, Nicholas Marshall informed us of the identity
$\|\sqrt{(k-1)^{2}+2}+\sqrt{(k+1)^{2}+2}+\sqrt{(2k)^{2}-8}\|\sim\frac{4}{k^{5}}$
which gives an infinite family of examples showing $\|\sqrt{a}+\sqrt{b}+\sqrt{c}\|\lesssim n^{-5/2}$. Roger Heath-Brown reports a similar identity, there appear to be many. All these examples have in common that each square root is a small perturbation of an integer (i.e. of the form $\sqrt{k^{2}+\mbox{small}}$). One would perhaps be inclined to believe that there should be many ‘non-perturbative’ examples that are $\ll n^{-5/2}?$ Tomás Oliveira e Silva informed me of the example
$\|\sqrt{11075}+\sqrt{27187}+\sqrt{68057}\|=1.26\cdot 10^{-15}\sim\frac{0.4}{68057^{3}}$
which is maybe suggestive of the bound $\lesssim n^{-3}$. Since we have $\sim n^{3}$ expressions mod 1, if they behaved randomly, the scaling would be $n^{-3}$. In that sense, one of way of posing the question is whether the sum of three roots is sufficiently ‘random’ to forget about the algebraic structure of each individual summand.
There are a number of interesting related problems, some of which might be easier. One that appears strictly easier concerns the set
$A=\left\{\sqrt{a}+\sqrt{b}\mod 1:1\leq a,b\leq n\right\}\subset[0,1].$
It might be interesting to have a better understanding of its gap structure. We note that $A\cap[0,c/n^{3/2}]=\emptyset$ which follows from Angluin-Eisenstat argument. Is it true that $A$ intersects every interval $J\subset[0,1]$ with length $|J|\gtrsim n^{-3/2}$ Is it true that there are ‘few’ large gaps at scale $\sim n^{-3/2}$?
@@
Remarks.
  1. (1)
    The proof uses both the van der Corput method (needing $k\geq 241$) and the Vinogradov method: the implicit constant $c>0$ is very small and
  2. (2)
    The argument presented in Proposition 1 generalizes to imply the lower bound $\gtrsim n^{1/2-2^{k-1}}$. A polynomial rate (in $k$) appears to be out of reach.
  3. (3)
    The proof shows a slightly stronger result: we show that the largest gap in the set $\sqrt{a_{1}}+\dots+\sqrt{a_{k}}\mod 1\subset[0,1]$ is $\leq c_{k}\cdot n^{-c\cdot k^{1/3}}$. In particular, while there are summands that are guaranteed to be close to an integer, the method also guarantees the existence of summands close to, say, $\mathbb{N}+\sqrt{2}$.
@@ 1.5. Exponential sum estimates
We will argue that for a suitable choice of $s\ll n^{k/2}$, the exponential sums are sufficiently small to be absorbed by the main term which then guarantees the existence of (many) solutions. By considering the modified function $h_{y}=h(x-y)$ and realizing that $\widehat{h_{y}(\ell)}$ is merely a phase-shift away from $\widehat{h}(\ell)$, a phase-shift that immediately disappears when applying the triangle inequality, we see that the argument can be equally applied to $h(x-y)$ and thus really is a statement about the number of solutions close to any arbitrary $0\leq y\leq 1$ (not just 0).
The main bottleneck in the argument appears to be getting good control on the exponential sum. The exponent pair hypothesis
@@ 2. Proof
One of the advantages is that the exponential sum $\sum_{1\leq a\leq n}e^{2\pi i\ell\sqrt{a}}$ is completely explicit. One way one could hope to avoid it is to relate it to the continuous integr to relate
@@
The presence of a completely explicit term that induces further oscillations does seem to inspire some hope that are more precise understanding might be possible, however, it does not appear to be straightforward.
Acknowledgment. The author is grateful for interesting discussions with Roger Heath-Brown, Arturas Dubickas, Charles Greathouse, Nicholas Marshall, Tomás Oliveira e Silva and Victor Reis. I am especially indebted to Igor Shparlinski for insightful remarks and finding a gap in an earlier version of the manuscript.